OpenStudy (anonymous):

A group of people were given a personality test to determine if they were Type A or Type B. The results are shown in the table below: Male Female Type A 55 75 Type B 48 22 Compare P(Female or Type B) with P(Female | Type B).

OpenStudy (anonymous):

a. There is not enough information b. P(Female or Type B) > P(Female | Type B) c. P(Female or Type B) = P(Female | Type B) d. P(Female or Type B) < P(Female | Type B) IS B CORRECT?

OpenStudy (anonymous):

I will award medal and fan

OpenStudy (kirbykirby):

\[P(\text{ female or type B})=P(\text{female})+P(\text{type B})-P(\text{female AND type B)} \\ \, \\ P(\text{female | Type B})=\frac{P(\text{female AND type B})}{P(\text{type B})}\]

OpenStudy (anonymous):

OR = 97 + 70 - 22 = 145 FEMALE | TYPE B = 22 / 70

OpenStudy (anonymous):

or = 145 / 200 = 0.725 f | type b = 0.314

OpenStudy (anonymous):

so its B??

OpenStudy (kirbykirby):

yes :)

OpenStudy (anonymous):

thank you!

OpenStudy (kirbykirby):

your welcome !